Find the vector equation of the plane determined by point ( 1,1,-1) (1, 1,1) and (0,1,1). Also find point of intersection of line r = (1+3t)I + (2-t) j + (1+t) and the plane.
1
Expert's answer
2019-03-14T13:03:40-0400


let A be(1,1,-1)

B (1,1,1)

C (0,1,1)

then "\vec{AB} = (0,0,2)" and "\vec{AC} = (-1,0,2)"

the vector equation will be ( Rm-Ra, AB, AC) = 0, where m (x,y,z)


coordinate form (matrix):

"\begin{vmatrix}
 x-1 & y-1 & z+1 \\
 0 & 0 & 2 \\
-1 & 0 & 2
\end{vmatrix}" = 0


"r= \vec{Ra}+S* \vec{AB} +t* \vec{AC}"

where S and t are some coefficients


vector-coordinate form:

"\begin{cases}
x= 1+(-1)*t \\
y=1 \\
z= -1 + 2*S + 2*t
\end{cases}" (1)


"line: \vec{r} = (1+3*t)* \vec{i} + (r-t)* \vec{j}+ (1+t)* \vec{k}"


"\begin{cases}
x=1+3t' \\
y=2-t' \\
z=1+t'
\end{cases}" (2)


from (1) y=1, then from (2) t' = 2-y = 2-1=1

then from (2) t'= z-1 follows 1=z-1, z=2

then from (2) x=1+3*1=4


answer:

point of intersection (4,1,2)